# Pareto-optimality of pulses for robust population transfer in a ladder-type qutrit

### v2

@@ IV Parameter Analysis
Figure 3: The parameter sensitivity of (a) $\max_{t}(p_{\rm{f}})$, (b) detuning robustness $\Delta_{\rm{rob}}$, and (c) amplitude robustness $\Omega_{\rm{rob}}$, where the vertical axis is the standard deviation $\sigma$ and the horizontal axis is the mean absolute effect $\mu^{*}$ found via the Morris method with a $5\times 10^{-3}\%$ perturbative decrease on each model parameter. The legend is shown in (d).
@@
To assess the sensitivity of the model parameters on each objective outcome, we employ a modified version of the Morris sensitivity test [51]. This method involves calculating two key metrics: the mean absolute effect $\mu^{*}$ and the standard deviation $\sigma$.
@@
The elementary effect for a given parameter is calculated by perturbing the parameter by a fixed percentage (in our case, a $5\times 10^{-3}\%$ decrease) while keeping all other parameters fixed. The elementary effect is then defined as the difference between the perturbed and unperturbed model output, divided by the perturbation step size ($-5\times 10^{-5}$). Unlike the conventional Morris method, we use the relative elementary effect, which is obtained by dividing each elementary effect by the unperturbed output. The mean absolute effect is the average of the absolute values of these relative elementary effects, providing a measure of the overall importance of each parameter in influencing the model output. Explicitly, the mean absolute effect is
@@
where $E_{i}$ is the relative elementary effect for the $i$-th sample of a parameter and $N$ is the total number of samples (175 for each of our parameters).
@@
where $\mu$ is the mean effect and the $N-1$ in the denominator accounts for the potential variability beyond the analyzed set. Unlike $\mu^{*}$, which is the mean of the absolute values of the relative elementary effects (Eq. 11), $\mu$ considers the raw values and retains their signs. The mean effect is a useful metric in its own right, as it indicates whether a parameter’s influence on the output is typically positive or negative, revealing the directionality of its impact. Moreover, the ratios of $\mu/\mu^{*}$, which assume values between $-1$ and $1$, inform us of the type of correlation of the parameters relative to the outputs. We have calculated these ratios for $\max_{t}(p_{\rm{f}})$, detuning robustness, and amplitude robustness for the case of a $5\times 10^{-3}\%$ perturbative decrease on each parameter. For instance, we see that for $\max_{t}(p_{\rm{f}})$, the amplitude $\Omega_{0}$ and envelope order $n$ parameters, each have a completely positive correlation with $\max_{t}(p_{\rm{f}})$ for all models, i.e., $\mu=\mu^{*}$, so the increase in each of these parameters, keeping all others fixed, yields a higher $\max_{t}(p_{\rm{f}})$. The same can be said for the detuning parameter $k_{1}$, except for models of type-4 detuning, i.e., SG4 and SECH4 models. For detuning and amplitude robustness, the ratios $\mu/\mu^{*}$ largely lie between $-1$ and $1$ for the models, which suggests a more complex correlation. On the other hand, the envelope parameters $\beta$ and $\epsilon$ are completely negatively correlated, i.e., $\mu=-\mu^{*}$, for all models when the output is $\max_{t}(p_{\rm{f}})$, and nearly completely negatively correlated when the output is detuning robustness. However, for amplitude robustness, these parameters are generally strongly positively correlated for the models.
@@
As can be seen in Fig. 3 (a), there is a cluster of $\Omega_{0}$ parameters with high $\mu^{*}$ and low $\sigma$. This cluster solely contains $\Omega_{0}$ for each model and pulse duration, which clearly indicates $\Omega_{0}$ has the most profound effect on the outcome of minimizing $\max_{t}(p_{\rm{f}})$. In Fig. 3 (b) and (c) the standard deviation is generally much higher among the parameters.
V Discussion and conclusions
@@
Inspired by exactly solvable two-state models, we have investigated the Pareto-optimality of population transfer from the ground state to the first excited state in a ladder-type qutrit. Our optimization considered two competing objectives: minimizing the maximum transient population of the second excited state, and maximizing detuning robustness. We define the latter as the width of the first excited state population $p_{\rm{e}}$ curve between the first two points where $p_{\rm{e}}(T/2)=0.99$, provided that $p_{\rm{e}}(T/2)\geq 0.999$ is reached at least once within this interval. Due to the inherent tradeoff between these objectives, the Pareto front represents the set of optimal solutions to this multiobjective problem.
To explore the optimal pulse design, we analyzed eight pulse models, each defined by a combination of an envelope function and a detuning function. The envelope function is either a super-Gaussian or a hyperbolic secant, while the detuning function follows one of four forms: linear (type-1), quintic (type-2), hyperbolic tangent (type-3), or a superposition of hyperbolic secant and hyperbolic tangent (type-4). Each model satisfies the conditions: $\Omega(0)=\Omega_{0}$, $\Omega(\pm T/2)=0$, and $\Delta(0)-\delta=0$. We further examined pulse durations of 50 ns and 200 ns and determined the Pareto-optimal model parameters that achieve the best tradeoff between our objectives.
As shown in Fig. 2, the resulting Pareto fronts cluster into two distinct regions, with the 200 ns pulses yielding more optimal results. To further contextualize these findings, we approximate theoretical bounds using the Landau-Zener formula. By expanding the nonlinear detuning functions around $t=0$ and extracting the linear term (Landau-Zener velocity), we compare our results against this theoretical benchmark. We find that the Pareto fronts—particularly those corresponding to type-3 and type-4 detunings—closely approach the Landau-Zener boundary. Additionally, our analysis reveals that the choice of detuning function has a more significant impact on optimality than choosing between the two flat-at-the-top envelope functions.
Further, all pulse models considered are capable of delivering moderate detuning robustness (e.g., $\Delta_{\rm{rob}}=0.1$ GHz) while keeping the maximum transient population of the second excited state low (i.e., $\max_{t}(p_{\rm{f}})$ $\leq 0.01$). However, achieving higher detuning robustness necessitates the use of nonlinear detuning profiles (types 2–4). Shorter pulses ($T=50$ ns) generally need higher $\max_{t}(p_{\rm{f}})$ to achieve similar robustness, while longer pulses ($T=200$ ns) can reach comparable or even better robustness with lower transient excitation.
Moreover, this trend that all models perform comparably well at moderate detuning robustness but diverge at higher values— is reflected in the structure of the Pareto fronts. For detuning robustness approximately up to 0.1 GHz (50 ns pulses) and 0.15 GHz (200 ns pulses), all models exhibit comparable $\max_{t}(p_{\rm{f}})$, as their fronts cluster tightly into two distinct groups corresponding to the pulse durations. This suggests that within these ranges, performance is largely model-agnostic.
To assess the sensitivity of each model parameter, we employed the Morris method, perturbing each parameter individually and evaluating its effect on $\max_{t}(p_{\rm{f}})$, $\Delta_{\rm{rob}}$, and $\Omega_{\rm{rob}}$. The relative elemental effect is computed as the difference between the perturbed and unperturbed outputs divided by the signed step size and further divided by the unperturbed output. From this, we compute the mean effect ($\mu$), mean absolute effect ($\mu^{*}$), and standard deviation ($\sigma$) of each parameter’s influence. By plotting $\sigma$ versus $\mu^{*}$ (Fig. 3), we identify key trends: For $\max_{t}(p_{\rm{f}})$, the amplitude $\Omega_{0}$ stands out as the most impactful parameter, forming an island of high $\mu^{*}$ and low $\sigma$ across all models. For $\Delta_{\rm{rob}}$ and $\Omega_{\rm{rob}}$, no single parameter dominates, suggesting a more intricate interplay between multiple parameters. However, some exceptions exist, such as $\Omega_{0}$, which has no impact on amplitude robustness.

### v4

@@ 4 Parameter Analysis
Figure 3: The parameter sensitivity of (a) $\max_{t}(p_{\rm{f}})$, (b) detuning robustness $\Delta_{\rm{rob}}$, and (c) amplitude robustness $\Omega_{\rm{rob}}$, where the vertical axis is the standard deviation $\sigma$ and the horizontal axis is the mean absolute effect $\mu^{*}$ found via the Morris method with a $5\times 10^{-3}\%$ perturbative decrease on each model parameter. The clustering of $k_{1}$ and $n$ is shown with higher resolution in (b) and (c), respectively, via the orange-colored insets. The legend is shown in (d).
@@
To assess the impact of the model parameters on each objective outcome, we employ a modified version of the Morris sensitivity test [Morris1991]. This method involves calculating two key metrics: the mean absolute effect $\mu^{*}$ and the standard deviation $\sigma$.
@@
The elementary effect for a given parameter is calculated by perturbing the parameter by a fixed percentage (in our case, a $5\times 10^{-3}\%$ decrease) while keeping all other parameters fixed. The elementary effect is then defined as the difference between the perturbed and unperturbed model output, divided by the perturbation step size. Unlike the conventional Morris method, we use the relative elementary effect, which is obtained by dividing each elementary effect by the unperturbed output. The mean absolute effect is the average of the absolute values of these relative elementary effects, providing a measure of the overall importance of each parameter in influencing the model output. Explicitly, the mean absolute effect is
@@
where $E_{i}$ is the relative elementary effect for the $i$-th sample of a parameter, and $N$ is the total number of samples (175 for each of our parameters).
@@
where $\mu$ is the mean effect and the N1N-1 in the denominator accounts for the potential variability beyond the analyzed set. Unlike $\mu^{*}$, which represents the mean of the absolute values of the relative elementary effects as defined in Eq. (11), $\mu$ considers the raw values and retains their signs. The mean effect is a useful metric in its own right, as it indicates whether a parameter’s influence on the output is typically positive or negative, revealing the directionality of its impact. Moreover, the ratios of $\mu/\mu^{*}$, which assume values between 1-1 and $1$, inform us of the type of correlation of the parameters relative to the outputs. In Appendix A, we have calculated these ratios for $\max_{t}(p_{\rm{f}})$, detuning robustness, and amplitude robustness for each parameter. For instance, we see that for $\max_{t}(p_{\rm{f}})$, the amplitude $\Omega_{0}$ and envelope order $n$ each have a completely positive correlation with $\max_{t}(p_{\rm{f}})$ for all models, i.e., $\mu=\mu^{*}$, so the increase in each of these parameters, keeping all others fixed, yields a higher $\max_{t}(p_{\rm{f}})$. The same can be said for the detuning parameter $k_{1}$, except for models of type-4 detuning, i.e., SG4 and SECH4 models. For detuning and amplitude robustness, the ratios $\mu/\mu^{*}$ largely lie between 1-1 and $1$ for the models, which suggests a more complex correlation. On the other hand, the envelope parameters $\beta$ and $\epsilon$ are completely negatively correlated, i.e., $\mu=-\mu^{*}$, for all models when the output is $\max_{t}(p_{\rm{f}})$ and nearly completely negatively correlated when the output is detuning robustness. However, for amplitude robustness, these parameters are generally strongly positively correlated for the models.
@@
As shown in Fig. 3(a), there is a clear cluster of $\Omega_{0}$ parameters with high mean absolute effect ($\mu^{*}$) and low standard deviation ($\sigma$) across all models and pulse durations. This indicates that $\Omega_{0}$ has the strongest and most consistent influence on minimizing the maximum transient second excited state population $\max_{t}(p_{\rm f})$.
In contrast, Fig. 3(b) and (c) display generally higher variability among parameters, indicating a more complex interplay. However, Fig. 3(b) reveals a prominent high-$\mu^{*}$ cluster for $k_{1}$, suggesting that $k_{1}$ plays a dominant role in determining detuning robustness $\Delta_{\rm rob}$—an aspect analyzed further in Appendix B. Amplitude robustness, in particular how it relates to detuning robustness, is discussed in Appendix C. We note in Fig. 3(c) that there is a noticeable cluster of envelope order $n$, suggesting it is an impactful parameter in determining amplitude robustness.
5 Discussion and conclusions
@@
Inspired by exactly solvable two-state models, we have investigated the Pareto-optimality of population transfer from the ground state to the first excited state in a ladder-type qutrit. Our optimization considered two competing objectives: minimizing the maximum transient population of the second excited state and maximizing detuning robustness. We define the latter as the width of the first excited state population $p_{\rm{e}}(T/2)$ curve between the first two points where $p_{\rm{e}}(T/2)=0.99$, provided that $p_{\rm{e}}(T/2)\geq 0.999$ is reached at least once within this interval. Due to the inherent trade-off between these objectives, the Pareto front represents the set of optimal solutions to this multiobjective problem.
We explored eight pulse models, combining two envelope functions (super-Gaussian or hyperbolic secant) with four detuning profiles: linear (type-1), quintic (type-2), hyperbolic tangent (type-3), and a superposition of hyperbolic secant and hyperbolic tangent (type-4). All pulses satisfy $\Omega(0)=\Omega_{0}$, $\Omega(\pm T/2)=0$, and $\Delta(0)-\delta=0$. We optimized parameters for two pulse durations, $T=50$ ns and $T=200$ ns.
The Pareto fronts, shown in Fig. 2(a), reveal two clusters based on pulse duration, with longer pulses yielding superior trade-offs. Comparing with Landau-Zener-based bounds derived from local linearization of detuning functions, we find that the $T=50$ ns pulses lie outside the adiabatic regime, especially at higher values of detuning robustness. Among the $T=200$ ns pulses, only the models with nonlinear detunings (types 2–4) lie inside the Landau-Zener bounds when delivering high robustness where the models with type-3 and type-4 detunings perform best. As such, our results suggest that detuning shape has a greater impact than envelope choice.
All pulse models perform similarly at moderate detuning robustness—up to about 0.1 GHz (50 ns) and 0.15 GHz (200 ns)—with low transient excitation (i.e., $\max_{t}(p_{\rm{f}})\leq 0.01$) in all cases, as shown by the Pareto fronts. However, achieving higher detuning robustness generally requires nonlinear detuning profiles (types 2–4). Longer pulses also tend to improve performance by further reducing transient excitation.
Finally, we applied the Morris method to assess parameter sensitivity. The most influential parameter for $\max_{t}(p_{\rm{f}})$ is the amplitude $\Omega_{0}$, consistently showing high mean absolute effect and low standard deviation. For detuning and amplitude robustness, influence is generally more distributed, with the exception of the detuning prefactor $k_{1}$, which clearly dominates for detuning robustness.
@@
[…]
Competing interests The authors declare no competing interests.
[…]
Ethics approval and consent to participate Not applicable.
[…]
Consent for publication Not applicable.
[…]
Data Availability Data is available upon reasonable request.
[…]
Materials availability Not applicable.
[…]
Code availability Codes are available upon reasonable request.
[…]
Author contribution J.J.M. performed the numerical experiments and data processing. J.J.M. and M.K. developed the methodology and performed the analysis. M.K. proposed the initial research idea and together with G.S.P. supervised the project. All authors contributed to writing of the manuscript.
Appendix A Parameter correlations with the outputs
Following the Morris method as described in Sect. 4, we have tabulated $\mu/\mu^{*}$ for outputs $\max_{t}(p_{\rm{f}})$, $\Delta_{\rm{rob}}$, and $\Omega_{\rm{rob}}$ in Tables A.1, A.2, A.3, respectively. These ratios necessarily range from 1-1 to $1$, where a ratio of 1-1 indicates a parameter is completely negatively correlated with the output, and a ratio of $1$ indicates a parameter is completely positively correlated with the output.
A.1 Parameter correlations with respect to $\max_{t}(p_{\rm{f}})$
We see in Table A.1 that for $\max_{t}(p_{\rm{f}})$, the ratios of $\mu/\mu^{*}$ corresponding to the envelope parameters $\Omega_{0}$ and $n$ are unity for each model and pulse duration, suggesting that they are completely positively correlated with output. However, for the other envelope parameter $\beta\backslash\epsilon$ the opposite is true; i.e., the ratios are all negatively correlated with $\max_{t}(p_{\rm{f}})$. Among the detuning parameters, $k_{1}$ and $\gamma_{1}$ are both completely positively correlated; whereas the results for the other parameters are more mixed, with the notable exception of $k_{5}$, where the ratios suggest strong negative correlations with the objective. It is also worth emphasizing that the parameter $k_{2}$ for the type-2 detuning models and the type-4 detuning functions are fundamentally different: in the former, $k_{2}$ controls the convexity or concavity of the detuning function and in the latter, $k_{2}$ determines the size of the dip which the hyperbolic secant term introduces in the detuning function. Thus, this discrepancy in the functionality of this parameter among the type-2 and type-4 detuning functions is reflected in the more extreme values of the ratios for the models with type-2 detuning compared to those of type-4, with the exception of the SECH4 ($T=200$ ns) model, where the ratios are interestingly 1-1. In order to meet our Pareto objective which involves the minimization of $\max_{t}(p_{\rm{f}})$ over $t\in[-T/2,T/2]$, Table A.1 informs us that we should avoid high values of $\Omega_{0}$, $n$, $k_{1}$, and $\gamma_{1}$, while seeking more extreme values, i.e., further from zero, of $\beta\backslash\epsilon$ and $k_{5}$.
Table A.1: Ratios of $\mu/\mu^{*}$ for $\max_{t}(p_{\rm{f}})$
Model ($T$) $\Omega_{0}$ $n$ $\beta\backslash\epsilon$ $k_{1}$ $k_{2}$ $k_{3}$ $k_{4}$ $k_{5}$ $\gamma_{1}$ $\gamma_{2}$
SG1 (50 ns) 1 1 -1 1 - - - - - -
SG1 (200 ns) 1 1 -1 1 - - - - - -
SECH1 (50 ns) 1 1 -1 1 - - - - - -
SECH1 (200 ns) 1 1 -1 1 - - - - - -
SG2 (50 ns) 1 1 -1 1 0.91 -0.81 -0.77 -0.93 - -
SG2 (200 ns) 1 1 -1 1 -0.99 -0.91 0.2 -0.99 - -
SECH2 (50 ns) 1 1 -1 1 -0.97 0.95 0.38 -0.91 - -
SECH2 (200 ns) 1 1 -1 1 -0.95 -0.36 -0.65 -0.94 - -
SG3 (50 ns) 1 1 -0.99 1 - - - - 1 -
SG3 (200 ns) 1 1 -1 1 - - - - 1 -
SECH3 (50 ns) 1 1 -1 1 - - - - 1 -
SECH3 (200 ns) 1 1 -1 1 - - - - 1 -
SG4 (50 ns) 1 1 -0.99 1 0.01 - - - 1 0.057
SG4 (200 ns) 1 1 -1 1 -0.46 - - - 1 -0.46
SECH4 (50 ns) 1 1 -1 1 0.047 - - - 1 0.057
SECH4 (200 ns) 1 1 -1 1 -1 - - - 1 -1
Table A.1: The ratios of $\mu/\mu^{*}$ for $\max_{t}(p_{\rm{f}})$ for each model and pulse duration when each parameter is independently perturbed by a $5\times 10^{-3}\%$ decrease following the Morris method. Positive ratios (red) indicate an overall positive correlation and negative ratios (blue) indicate an overall negative correlation.
A.2 Parameter correlations with respect to $\Delta_{\rm{rob}}$
The $\mu/\mu^{*}$ ratios for detuning robustness $\Delta_{\rm{rob}}$ are tabulated in Table A.2 for each of the models. For the envelope parameters, we again see the same trend as in Table A.1, though with generally less extreme values. We note, for instance, that the ratios for $\Omega_{0}$ are all such that $0.5\leq\mu/\mu^{*}\leq 0.97$. For the parameter $n$, the ratios generally suggest a stronger positive correlation than for $\Omega_{0}$, and the ratios for $\beta\backslash\epsilon$ are all negative, but with less extreme values than in Table A.1. Further, we note that the ratios corresponding to $n$ and $\beta\backslash\epsilon$ appear to be less extreme for the longer pulse duration, suggesting that a longer pulse duration for these parameters yields inconsistent directional effects. The ratios corresponding to each of the detuning parameters are also less extreme in Table A.2, with the $\mu/\mu^{*}$ ratios of $k_{1}$ and $\gamma_{1}$ suggesting strong positive correlation with $\Delta_{\rm{rob}}$. For the other parameters, the ratios generally suggest that the models have inconsistent effects relative to each other.
Table A.2: Ratios of $\mu/\mu^{*}$ for $\Delta_{\rm{rob}}$
Model ($T$) $\Omega_{0}$ $n$ $\beta\backslash\epsilon$ $k_{1}$ $k_{2}$ $k_{3}$ $k_{4}$ $k_{5}$ $\gamma_{1}$ $\gamma_{2}$
SG1 (50 ns) 0.76 1 -1 0.98 - - - - - -
SG1 (200 ns) 0.7 0.85 -0.82 0.53 - - - - - -
SECH1 (50 ns) 0.95 1 -1 0.93 - - - - - -
SECH1 (200 ns) 0.69 0.61 -0.63 0.5 - - - - - -
SG2 (50 ns) 0.88 1 -1 0.99 0.57 -0.56 -0.82 0.13 - -
SG2 (200 ns) 0.93 0.94 -0.83 0.83 -0.4 -0.41 0.045 0.084 - -
SECH2 (50 ns) 0.97 0.99 -1 0.98 -0.99 -0.75 0.55 -0.44 - -
SECH2 (200 ns) 0.71 0.94 -0.96 0.96 -0.7 0.55 -0.047 -0.28 - -
SG3 (50 ns) 0.6 0.95 -1 0.99 - - - - 1 -
SG3 (200 ns) 0.5 0.92 -0.88 0.84 - - - - 0.99 -
SECH3 (50 ns) 0.91 0.97 -0.99 0.97 - - - - 1 -
SECH3 (200 ns) 0.69 0.91 -0.96 0.92 - - - - 0.97 -
SG4 (50 ns) 0.81 0.96 -1 1 -0.81 - - - 0.99 -0.81
SG4 (200 ns) 0.79 0.97 -0.97 0.92 -0.5 - - - 0.97 -0.7
SECH4 (50 ns) 0.81 0.98 -1 0.99 -0.38 - - - 0.99 -0.31
SECH4 (200 ns) 0.55 0.71 -0.75 0.76 -0.38 - - - 0.78 0.55
Table A.2: The ratios of $\mu/\mu^{*}$ for $\Delta_{\rm{rob}}$ for each model and pulse duration when each parameter is independently perturbed by a $5\times 10^{-3}\%$ decrease following the Morris method. Positive ratios (red) indicate an overall positive correlation and negative ratios (blue) indicate an overall negative correlation.
A.3 Parameter correlations with respect to $\Omega_{\rm{rob}}$
Table A.3 shows the $\mu/\mu^{*}$ ratios for amplitude robustness $\Omega_{\rm{rob}}$. The ratios suggest generally weak and mixed directional effects for the amplitude parameter $\Omega_{0}$, which is not at all surprising as $\Omega_{0}$ principally has no effect on amplitude robustness. Further, we observe the ratios for the parameters $n$ and $\beta\backslash\epsilon$ indicate that they are generally strongly negatively correlated and positively correlated with $\Omega_{\rm{rob}}$, respectively. This correlation is opposite to what was observed for $\max_{t}(p_{\rm{f}})$ and $\Delta_{\rm{rob}}$. Across most models, $\mu/\mu^{*}$ of $k_{1}$ suggests negative correlation with amplitude robustness to varying degrees. The exceptions are for the SG4 ($200$ ns) and SECH4 ($200$ ns) models, where $k_{1}$ has the effect of scaling the detuning function, rather than determining its slope as in models of type-1 and type-2 detuning functions. Finally, we note that the ratios corresponding to $\gamma_{1}$ are indicative of a negative effect on $\Omega_{\rm{rob}}$. For the other parameters, the ratios generally suggest that the models have inconsistent effects relative to each other.
Table A.3: Ratios of $\mu/\mu^{*}$ for $\Omega_{\rm{rob}}$
Model ($T$) $\Omega_{0}$ $n$ $\beta\backslash\epsilon$ $k_{1}$ $k_{2}$ $k_{3}$ $k_{4}$ $k_{5}$ $\gamma_{1}$ $\gamma_{2}$
SG1 (50 ns) 0.2 -1 0.99 -0.79 - - - - - -
SG1 (200 ns) 0.42 -1 0.98 -0.81 - - - - - -
SECH1 (50 ns) -0.19 -1 0.97 -0.91 - - - - - -
SECH1 (200 ns) -0.087 -1 1 -0.99 - - - - - -
SG2 (50 ns) -0.044 -1 0.99 -0.47 -0.23 0.18 -0.57 -0.44 - -
SG2 (200 ns) -0.1 -1 0.99 -0.89 -0.44 -0.21 0.044 -0.21 - -
SECH2 (50 ns) 0.2 -0.98 0.9 -0.67 -0.76 0.72 -0.074 0.65 - -
SECH2 (200 ns) 0.23 -1 1 -0.87 -0.58 -0.3 -0.83 0.14 - -
SG3 (50 ns) -0.08 -0.99 0.99 -0.67 - - - - -0.79 -
SG3 (200 ns) 0.85 -1 1 -0.74 - - - - -0.98 -
SECH3 (50 ns) 0.067 -1 1 -0.43 - - - - -0.56 -
SECH3 (200 ns) -0.23 -1 1 -0.97 - - - - -0.99 -
SG4 (50 ns) 0.14 -0.99 1 -0.56 -0.89 - - - -0.78 -0.87
SG4 (200 ns) -0.052 -0.96 0.89 0.9 0.33 - - - -0.97 0.3
SECH4 (50 ns) 0.28 -0.99 0.99 -0.47 -0.82 - - - -0.61 -0.78
SECH4 (200 ns) -0.25 0.13 1 0.67 -1 - - - -0.32 -0.98
Table A.3: The ratios of $\mu/\mu^{*}$ for $\Omega_{\rm{rob}}$ for each model and pulse duration when each parameter is independently perturbed by a $5\times 10^{-3}\%$ decrease following the Morris method. Positive ratios (red) indicate an overall positive correlation and negative ratios (blue) indicate an overall negative correlation.
Appendix B Parameter Effects on the Pareto Objectives
This Appendix analyzes how the Pareto-optimal amplitude $\Omega_{0}$ and detuning prefactor $k_{1}$ influence the two Pareto objectives: the maximum transient second excited state population, $\max_{t}(p_{\rm{f}})$, which we aim to minimize, and the detuning robustness, $\Delta_{\rm rob}$, which we seek to maximize.
B.1 Effect of $\Omega_{0}$ on the Pareto Objectives
As observed in Sect. 3, longer pulses offer greater robustness to detuning. This explains the distinct clustering seen in Fig. B.1(a), where the Pareto-optimal amplitude is plotted against detuning robustness for both the 50 ns and 200 ns pulse durations of each model. Notably, the 50 ns pulses exhibit higher values of $\Omega_{0}$ than the 200 ns pulses at the same level of detuning robustness.
In Fig. B.1(b), the Pareto-optimal amplitudes $\Omega_{0}$ for each model and pulse duration approach zero as $\max_{t}(p_{\rm{f}})\rightarrow 0$. As $\max_{t}(p_{\rm{f}})$ increases, the amplitudes corresponding to shorter pulses ($T=50$ ns) tend to be higher than those for longer pulses ($T=200$ ns). Overall, a strong correlation is evident: $\Omega_{0}$ increases nearly monotonically with $\max_{t}(p_{\rm{f}})$, forming a universal curve that is independent of the pulse shape, duration, or detuning function. This trend is corroborated by the Morris analysis in Fig. 3(a), where $\Omega_{0}$ shows a high impact—characterized by a large mean absolute effect and low standard deviation—relative to the other parameters.
Figure B.1: The Pareto-optimal amplitudes $\Omega_{0}$ vs detuning robustness in panel (a) and $\max_{t}(p_{\rm{f}})$ in panel (b). Each model is evaluated at pulse durations $T=50$ ns (solid lines) and $T=200$ ns (dashed lines). The vertical arrows in (a) indicate that there are jumps in $\Omega_{0}$ corresponding to a change in sign of $k_{1}$.
B.2 Effect of $k_{1}$ on the Pareto objectives
Here, we investigate how different detuning profiles influence the Pareto objectives, with particular focus on the parameter $k_{1}$. As demonstrated by the Morris analysis in the main text, $k_{1}$ has the most significant impact. Moreover, as will be shown shortly, the detuning profiles are approximately linear—especially near the resonance point $\Delta(t)=0$—with higher-order terms contributing only minor corrections (cf. Fig. B.3).
As in the previous section, we begin by plotting the $k_{1}$ parameter as a function of the Pareto objectives, shown in Fig. B.2. There we observe that, opposite of what was found for $\Omega_{0}$, $k_{1}$ is highly correlated with the detuning robustness, and much less so with $\max_{t}(p_{\rm{f}})$. Namely, as a function of the detuning robustness, $k_{1}$ traces cluster into two universal curves, one for each of the pulse durations. Moreover, we find that $|k_{1}|$ is an approximately linear function of $\Delta_{\rm rob}$, with a sign switch occurring mid-way. The same discontinuity can be observed in the Pareto fronts: cf. Fig. 2(a) of the main text or Fig. B.1 where the discontinuities are marked with arrows.
To understand the origin of the discontinuity it is instructive to plot the detuning profiles on top of the qutrit spectrum. This is shown in Fig. B.3 for several different levels of detuning robustness. We observe that for small values of detuning robustness $k_{1}>0$, as we approach the $\mathrm{g-e}$ transition frequency from below (panel a). However, for larger values of $k_{1}$, $\Delta(t)$ will become nearly resonant with the two-photon $\mathrm{g-f}$ transition close to the beginning of the pulse which will inherently populate the second excited state. To avoid this, the sign of the detuning changes (see panels b and c), and the 2-photon resonance is driven only after the $\mathrm{g-e}$ transfer is complete, thus having little effect. To complete the analysis, we evaluate the Pareto front with a sign change of the detuning function $\Delta(t)\rightarrow-\Delta(t)$, from which we find that for small values of $k_{1}$$\Delta_{\rm{rob}}$ this leads to an increase of $\max_{t}(p_{\rm{f}})$; thus requiring a sign change of $k_{1}$ to achieve the optimal performance. This is illustrated in Fig. B.2(c) for the SG1 model with a pulse duration of $T=200$ ns, which is represented as a blue dotted line.
Fig. B.4 shows the $p_{\rm e}(T/2)$ vs $\delta$ traces for different $k_{1}$ and pulse durations. For an adiabatic trajectory, the transfer probability is set only by the Landau-Zener velocity $v=\left.\frac{d\Delta(t)}{dt}\right|_{t=t_{0}}$ and the spectral gap $\propto\Omega(t_{0})$ at the resonance condition $\Delta(t_{0})+\delta=0$. Since $\Delta(t)$ and $\Omega(t)$ are smooth functions of $t$, we expect $p_{\rm e}(T/2,\delta)$ to vary smoothly with respect to $\delta$. Instead, Fig. B.4 reveals the presence of finer structure, which we attribute to non-adiabatic trajectories, where the full quantum dynamics set the outcome. Naturally, preserving adiabaticity is harder for a short pulse duration, leading to more fluctuations in the transfer probability as a function of $\delta$. This, in turn, decreases the total detuning robustness for short pulses. In contrast, the $T=200$ ns traces show little to no structure, implying an adiabatic population transfer.
Figure B.2: The Pareto-optimal detuning prefactors $k_{1}$ as a function of (a) $\max_{t}(p_{\rm{f}})$ and (b) detuning robustness for each model at pulse durations $T=50$ ns (solid lines) and $T=200$ ns (dashed lines). Panel (c) shows the Pareto front (PF) with our usual definition of $\Delta(t)$ (solid blue line) and the Pareto front with $\Delta(t)\rightarrow-\Delta(t)$ (blue dotted line) for the SG1 model with a pulse duration of $T=200$ ns. The right vertical axis shows the values of the corresponding Pareto-optimal $k_{1}$ (solid red line).
Figure B.3: The Pareto-optimal detuning functions for each model at pulse durations $T=50$ ns (solid lines) and $T=200$ ns (dashed lines). Panel (a) shows the Pareto-optimal $\Delta(t)$ at the last positive $k_{1}$ index, whereas (b) shows the Pareto-optimal $\Delta(t)$ at the first negative $k_{1}$ index. Panel (c) shows the most detuning robust Pareto-optimal detuning function and (d) contains the legend.
Figure B.4: Plots showing detuning sweeps for each of the models using the Pareto-optimal parameters. Panels (a) and (b) show detuning sweeps with a robustness of approximately $\Delta_{\rm rob}=20$ MHz, for pulses of duration $T=50$ ns and $T=200$ ns, respectively. Panels (c) and (d) show the robustness of the models at the last $k_{1}$ index before $k_{1}<0$, for $T=50$ ns and $T=200$ ns, respectively. Panels (e) and (f) display the sweeps with maximum detuning robustness for $T=50$ ns and $T=200$ ns, respectively. For reference, a black dashed line is plotted at $p_{\rm e}(T/2)=0.99$ to indicate where the robustness (width) is measured.
Appendix C Amplitude Robustness
Although amplitude robustness was not one of the optimization objectives, we used the solutions of the Pareto-optimization to compute it. Thus, the amplitude robustness naturally satisfies the same criterion used to define detuning robustness—namely, the width of the $p_{\rm{e}}(T/2)$ curve between the first two points where $p_{\rm{e}}(T/2)=0.99$, provided that $p_{\rm{e}}(T/2)\geq 0.999$ is reached at least once within this interval.
Fig. C.1 shows a log-log scale of amplitude robustness versus detuning robustness for each model at the two pulse durations $T=50$ ns (a) and $T=200$ ns (b). The approximate linearity of the models and their convergence at higher robustnesses demonstrate that amplitude robustness and detuning robustness are highly correlated. Further, we see that the amplitude robustness $\Omega_{\rm rob}$ for the $T=50$ ns pulses and the $T=200$ ns pulses are each well-described by a linear relationship with respect to detuning robustness, i.e., $\Omega_{\rm rob}=1.1\Delta_{\rm rob}$ and $\Omega_{\rm rob}=1.2\Delta_{\rm rob}$, respectively.
Finally, in Fig. C.2 we show the amplitude robustness for each model by plotting $p_{\rm e}(T/2)$ as a function of the amplitudes $\Omega_{0}$, for which we use a logarithmic scale to provide a better resolution. In particular, Fig. C.2(a) and (b) show amplitude sweeps with effectively no robustness, i.e., $\Omega_{\rm rob}\approx 20$ MHz for $T=50$ ns and $T=200$ ns pulses, respectively, and Fig. C.2(c) and (d) show the traces with maximum amplitude robustness. Note that Fig. C.2(d) which is for $T=200$ ns pulses, only shows the full amplitude robustness for the models with linear detuning, as the nonlinear detuning models have greater amplitude robustness.
Figure C.1: A log-log scale plot showing amplitude robustness $\Omega_{\rm rob}$ vs detuning robustness $\Delta_{\rm rob}$ for each model evaluated at pulse durations (a) $T=50$ ns and (b) $T=200$ ns. The orange reference lines are the best-fit lines, each of which has a power of 1 and a slope near unity, indicating a near one-to-one relationship between $\Delta_{\rm rob}$ and $\Omega_{\rm rob}$.
Figure C.2: Semi-log scale plots showing amplitude sweeps for each of the models using the Pareto-optimal parameters. Panels (a) and (b) show amplitude sweeps with a robustness of approximately $\Omega_{\rm rob}=20$ MHz for pulses of duration $T=50$ ns and $T=200$ ns, respectively. Panels (c) and (d) display the sweeps with maximum amplitude robustness for $T=50$ ns and $T=200$ ns, respectively. For reference, a black dashed line is plotted at $p_{\rm e}(T/2)=0.99$ to indicate where the robustness (width) is measured.